# Globalization of perturbative Chern-Simons theory on the moduli space of flat connections in the BV formalism

### v2

@@ 4.3.1. Total horizontality (modulo $\Delta$-exact terms) on $\mathrm{FC}\times\mathrm{FC}$
(133) $\nabla^{\mathrm{tot}}Z=i\hbar\Delta_{\mathsf{a}}\Big(r_{{\widetilde{\delta A^{\prime}}}}(\mathsf{a})Z_{A,A^{\prime}}(\mathsf{a})\Big).$
@@ 4.4. Extension of $Z_{A,A^{\prime}}$ to a horizontal nonhomogeneous form in $A^{\prime}$
Consider the Feynman graph part of $\widehat{Z}$, $\sum_{\Gamma}\widehat{\Phi}_{\Gamma}$, with $\Gamma$ running over possibly disconnected graphs, with edges decorated by $\widehat{K}$ and leaves decorated by $\widehat{i}(\mathsf{a})$. (We include powers of $\hbar$ and the symmetry factor in $\widehat{\Phi}$.) The action of $\delta_{A^{\prime}}$ on $\widehat{\Phi}_{\Gamma}$ can be computed as a sum of (a) decorations of one edge of $\Gamma$ with $(\delta_{A^{\prime}}+[d_{A},-])\widehat{K}=1-\widehat{i}\,\widehat{p}$ (140a), plus (b) decorations of one leaf of $\Gamma$ with $(\delta_{A^{\prime}}+d_{A})\widehat{i}=\widehat{i}\widehat{\Theta}$ (140b). Upon summing over graphs, contributions of $1$ cancel out; contributions of $\widehat{i}\widehat{p}$ yield $i\hbar\Delta_{\mathsf{a}}\sum_{\Gamma}\widehat{\Phi}_{\Gamma}$; contributions of $\widehat{i}\widehat{\Theta}$ add up to
@@ 4.5. Metric dependence of the desynchronized partition function
(154) $\delta_{g}Z_{A,A^{\prime}}^{\mathrm{ren}}(\mathsf{a})=i\hbar\Delta_{\mathsf{a}}\left(r_{\delta g}(\mathsf{a})Z^{\mathrm{ren}}_{A,A^{\prime}}(\mathsf{a})\right)$
@@
This proof is analogous to the proof of Proposition 4.11, using the fact that the renormalization cancels potential contributions from hidden boundary strata. ∎
@@
Again, we note that since $K$ reduces the form degree along $M$, these sums are finite: The sum stops at $k=2$ for $\widehat{i}_{\delta g}$, $\widehat{p}_{\delta g}$ and $\widehat{K}_{\lambda_{\delta g}}$.
@@ 5.2. A formula for the Grothendieck connection
(194) $\Xi=-\langle[\delta A],\mathsf{a}\rangle+\frac{1}{2}\langle\mathsf{a},\check{\Theta}(\mathsf{a})\rangle+\sum_{T}\frac{1}{|\mathrm{Aut}(T)|}\check{\Phi}_{T}(\mathsf{a})\Big|_{\Omega^{1}(\underline{\mathcal{U}})}\quad\in\Omega^{1}(\underline{\mathcal{U}},\widehat{\mathrm{Sym}}(H_{A}[1])^{*})$
with $T$ running over trivalent trees, with notations as in (171). $\Xi$ splits as a 1-from along $A$ plus a 1-from along $A^{\prime}$ plus a 1-form along $g$:
@@
We will write $\widetilde{\mathsf{a}}=B(\mathsf{a}+\epsilon\mathsf{b})$ (all computations are $\bmod\,\epsilon^{2}$), with $B=\mathfrak{B}^{1}_{\widetilde{A}\leftarrow A;\widetilde{A}}\colon H^{1}_{A}\rightarrow H^{1}_{\widetilde{A}}$ the cohomology comparison map. [^52] Note that in the first order in $\epsilon$ it does not matter whether we choose $A^{\prime}=A$ or $A^{\prime}=\widetilde{A}$ in the cohomology comparison map $\mathfrak{B}_{\widetilde{A}\leftarrow A;A^{\prime}}$. Equation (202) reads
[…]
@@
(207) $=[\varphi_{A,A}(\mathsf{a})+\epsilon(d_{\mathsf{a}}\varphi_{A,A})(\alpha+\mathsf{b}+\beta)]$
@@
(210) $\det(B^{\vee})\circ\underline{Z}_{\underline{\varphi}_{A}(\alpha)}(B(\mathsf{a}))=\underline{Z}_{A}(\alpha+\mathsf{a})+i\hbar\Delta_{\mathsf{a}}R(A,\alpha,\mathsf{a}),$
@@
(213) $\nabla^{G}\underline{Z}=i\hbar\Delta_{\mathsf{a}}(\underline{r}\,\underline{Z}),$
@@
(214) $\nabla^{G,\mathrm{triv}}\underline{Z}=i\hbar\Delta_{\mathsf{a}}(\pi_{*}I^{*}(r)\,\underline{Z}),$
@@
(240) $Z^{\mathrm{glob}}$ $=\sum_{k=0}^{\dim\mathcal{M}^{\prime}}\frac{(i\hbar)^{k}}{k!}\left\langle\frac{\partial}{\partial\mathsf{a}^{2}},\frac{\partial}{\partial[\delta A]}\right\rangle^{k}\widetilde{Z}^{(k)}\Big|_{\mathsf{a}^{1}=0}$
$=\left(e^{i\hbar\left\langle\frac{\partial}{\partial\mathsf{a}^{2}},\frac{\partial}{\partial[\delta A]}\right\rangle}\widetilde{Z}\right)\Big|_{\mathsf{a}^{1}=[\delta A]=0}.$
@@
(242) $\mathsf{K}\omega|_{\mathsf{a}^{1}\rightarrow 0}\sim\frac{1}{\deg\omega}\left\langle\mathsf{a}^{1},\frac{\partial}{\partial[\delta A]}\right\rangle\omega|_{\mathsf{a}^{1}=0},$
@@
Formula (240) can be seen as the perturbative evaluation of the following path integral:
(244) $Z^{\mathrm{glob}}(A)=\mathrm{RT}\int_{H^{2}_{A}[-1]\oplus H^{1}_{A}[1]}\mathcal{D}\mathsf{a}^{2}\,\mathcal{D}\zeta\int_{\mathcal{L}=\Omega_{d_{A}^{*}\mathrm{-ex}}[1]}\mathcal{D}\alpha_{\mathrm{fl}}\\ \exp\frac{i}{\hbar}\Bigg(S_{CS}(A+i(\mathsf{a}^{2})+\alpha_{\mathrm{fl}})+\langle\mathsf{a}^{2},\zeta\rangle\\ +\int_{M}\frac{1}{2}\langle\alpha_{\mathrm{fl}},d_{A}G\mathrm{ad}^{*}_{i(\zeta)}\alpha_{\mathrm{fl}}\rangle+\langle\alpha_{\mathrm{fl}},d_{A}G\mathrm{ad}^{*}_{i(\zeta)}i(\mathsf{a}^{2})\rangle\Bigg).$
@@
The symbol $\mathrm{RT}$ “remove trees” in (244) stands for removing from the perturbative expansion Feynman graphs with the property that if one cuts all dashed edges ($\mathsf{a}^{2}-\zeta$ propagators), one of the remaining connected components is a tree with exactly one $H$-vertex.
@@ 5.5. Metric dependence of the global partition function
(251) $\check{Z}^{\mathrm{glob,ren}}$ $=\sum_{k=0}^{\dim\mathcal{M}^{\prime}}\frac{(i\hbar)^{k}}{k!}\left\langle\frac{\partial}{\partial\mathsf{a}^{2}},\frac{\partial}{\partial[\delta A]}\right\rangle^{k}\overline{Z}^{\mathrm{ren}(\bullet,k)}\Big|_{\mathsf{a}^{1}=0}$
$=\left(e^{i\hbar\left\langle\frac{\partial}{\partial\mathsf{a}^{2}},\frac{\partial}{\partial[\delta A]}\right\rangle}\overline{Z}^{\mathrm{ren}}\right)\Big|_{\mathsf{a}^{1}=[\delta A]=0}.$
@@
(252) $(\delta_{g}+\nabla^{G}+\{\mu,-\}+\Delta_{\mathsf{a}}\mu-i\hbar\Delta_{\mathsf{a}})=0.$
@@
    The induced differential in (254) is $\sum_{k\geq 0}\mathsf{p}\varkappa(-\mathsf{K}\varkappa)^{k}\mathsf{i}$. Since the image of $\mathsf{i}$ is in zero-forms on $\mathcal{M}^{\prime}$, $K$ reduces the form degree on $\mathcal{M}^{\prime}$ by one and $\varkappa$ does not decrease the form degree on $\mathcal{M}^{\prime}$, all terms with $k>0$ vanish and the induced differential is $\mathsf{p}\varkappa\mathsf{i}=\delta_{g}-i\hbar\Delta$.
@@
(256) $\check{Z}^{\mathrm{glob,ren}}(A,\mathsf{a}^{2})=e^{\frac{i}{\hbar}c(\hbar)\frac{S_{\mathrm{grav}}(g,\phi)}{2\pi}}\cdot\mathrm{RT}\int_{H^{2}_{A}[-1]\oplus H^{1}_{A}[1]}\mathcal{D}\mathsf{a}^{2}\,\mathcal{D}\zeta\int_{\mathcal{L}=\Omega_{d_{A}^{*}\mathrm{-ex}}[1]}\mathcal{D}\alpha_{\mathrm{fl}}\\ \exp\frac{i}{\hbar}\Big(S_{CS}(A+i(\mathsf{a}^{2})+\alpha_{\mathrm{fl}})+\langle\mathsf{a}^{2},\zeta\rangle\\ +\int_{M}\underbrace{\frac{1}{2}\langle\alpha_{\mathrm{fl}},d_{A}G\mathrm{ad}^{*}_{i(\zeta)}\alpha_{\mathrm{fl}}\rangle+\langle\alpha_{\mathrm{fl}},d_{A}G\mathrm{ad}^{*}_{i(\zeta)}i(\mathsf{a}^{2})\rangle}_{-\frac{1}{2}\langle i(\mathsf{a}^{2})+\alpha_{\mathrm{fl}},H_{\delta A^{\prime}=i(\zeta)}(i(\mathsf{a}^{2})+\alpha_{\mathrm{fl}})\rangle}\underbrace{-\frac{1}{2}\langle\alpha_{\mathrm{fl}},\lambda_{\delta g}\alpha_{\mathrm{fl}}\rangle-\langle\alpha_{\mathrm{fl}},\lambda_{\delta g}i(\mathsf{a}^{2})\rangle}_{-\frac{1}{2}\langle i(\mathsf{a}^{2})+\alpha_{\mathrm{fl}},H_{\delta g}(i(\mathsf{a}^{2})+\alpha_{\mathrm{fl}})\rangle}\Big)$
with $H_{\delta A^{\prime}}$ and $H_{\delta g}$ as in (162b), (162c); $\mathsf{a}^{2}$ is interpreted as a vector in $T^{*}_{A}[-1]\mathcal{M}^{\prime}$.
@@
Here the graphical conventions are as in Figure 13; loose half-edges are decorated by $i(\mathsf{a}^{2})$; black circle vertex is decorated by $H_{\delta g}$. The graphs shown contribute in zeroth order in $\hbar$ and $\cdots$ is of order $\geq 1$ in $\hbar$.
The symbol $\mathrm{RT}$ in (256) stands for removing Feynman graphs such that if all dashed edges are deleted, there is a connected component which is a tree with one white vertex or one black vertex.
@@ Appendix D Some technical proofs
D.1. Proof of Proposition 2.9
@@
(317) $(d\widetilde{\kappa}_{A_{0}})_{\delta}=\mathrm{Id}+K_{A_{0}}\mathrm{ad}_{\delta}\colon\Omega^{1}\to\Omega^{1}$
is invertible. In particular, by the triangle inequality this happens when the operator norm of $K_{A_{0}}\delta$ is less than one.
@@
If $d_{A_{0}}\alpha=-\frac{1}{2}[\alpha,\alpha]$, then the first two terms cancel and the latter two terms vanish because $\alpha,\alpha]=-2d_{A_{0}}\alpha$ is exact. ∎
@@
However, by the results of Axelrod-Singer [AS91], the singular terms of the heat kernel expansion are proportional to $\mathrm{id}\in\mathrm{End}(\mathfrak{g})$ and hence vanish under the trace with $\mathrm{ad}_{\dot{A}_{0}}$ by unimodularity of $\mathfrak{g}$. Therefore, the zeta-regularized supertrace coincides with the point-splitting regularized supertrace that we use to define tadpoles in Feynman diagrams, cf. footnote 38.

### v3

@@ 4.3.1. Total horizontality (modulo $\Delta$-exact terms) on $\mathrm{FC}\times\mathrm{FC}$
(150) $\nabla^{\mathrm{tot}}Z=-i\hbar\Delta_{\mathsf{a}}\Big(r_{{\widetilde{\delta A^{\prime}}}}(\mathsf{a})Z_{A,A^{\prime}}(\mathsf{a})\Big).$
@@ 4.4. Extension of $Z_{A,A^{\prime}}$ to a horizontal nonhomogeneous form in $A^{\prime}$
Consider the Feynman graph part of $\widehat{Z}$, $\sum_{\Gamma}\widehat{\Phi}_{\Gamma}$, with $\Gamma$ running over possibly disconnected graphs, with edges decorated by $\widehat{K}$ and leaves decorated by $\widehat{i}(\mathsf{a})$. (We include powers of $\hbar$ and the symmetry factor in $\widehat{\Phi}$.) The action of $\delta_{A^{\prime}}$ on $\widehat{\Phi}_{\Gamma}$ can be computed as a sum of (a) decorations of one edge of $\Gamma$ with $(\delta_{A^{\prime}}+[d_{A},-])\widehat{K}=1-\widehat{i}\,\widehat{p}$ (157a), plus (b) decorations of one leaf of $\Gamma$ with $(\delta_{A^{\prime}}+d_{A})\widehat{i}=\widehat{i}\widehat{\Theta}$ (157b). Upon summing over graphs, contributions of $1$ cancel out; [^52] Cf. cancellation of principal boundary strata and footnote 49 in the proof of Proposition 4.11. contributions of $\widehat{i}\widehat{p}$ yield $i\hbar\Delta_{\mathsf{a}}\sum_{\Gamma}\widehat{\Phi}_{\Gamma}$; contributions of $\widehat{i}\widehat{\Theta}$ add up to
@@ 4.5. Metric dependence of the desynchronized partition function
(171) $\delta_{g}Z_{A,A^{\prime}}^{\mathrm{ren}}(\mathsf{a})=-i\hbar\Delta_{\mathsf{a}}\left(r_{\delta g}(\mathsf{a})Z^{\mathrm{ren}}_{A,A^{\prime}}(\mathsf{a})\right)$
@@
This proof is analogous to the proof of Proposition 4.11, using the fact that the renormalization cancels potential contributions from hidden boundary strata. [^55] By Corollary D.8, the restriction of the propagator to the boundary of configuration space in the desynchronized theory agrees modulo regular terms to the propagator in the case $A=A^{\prime}$, so the analysis of the hidden boundary strata of [AS94] carries over without change.
@@
Here $\mathrm{Met}$ denotes the Fréchet manifold of metrics and we are using its exterior calculus. Alternatively, one can restrict to an arbitrary finite-dimensional submanifold thereof. Again, we note that since $K$ reduces the form degree along $M$, the sums in (172) are finite: The sum stops at $k=2$ for $\widehat{i}_{\delta g}$, $\widehat{p}_{\delta g}$ and $\widehat{K}_{\lambda_{\delta g}}$.
@@ 5.2. A formula for the Grothendieck connection
(211) $\Xi=-\langle[\delta A],\mathsf{a}\rangle+\sum_{T}\frac{1}{|\mathrm{Aut}(T)|}\check{\Phi}_{T}(\mathsf{a})\Big|_{\Omega^{1}(\underline{\mathcal{U}})}\quad\in\Omega^{1}(\underline{\mathcal{U}},\widehat{\mathrm{Sym}}(H_{A}[1])^{*})$
with $T$ running over trivalent trees, with notations as in (188). $\Xi$ splits as a 1-form along $A$ plus a 1-form along $A^{\prime}$ plus a 1-form along $g$:
@@
We will write $\widetilde{\mathsf{a}}=B(\mathsf{a}+\epsilon\mathsf{b})$ (all computations are $\bmod\,\epsilon^{2}$), with $B=\mathfrak{B}^{1}_{\widetilde{A}\leftarrow A;\widetilde{A}}\colon H^{1}_{A}\rightarrow H^{1}_{\widetilde{A}}$ the cohomology comparison map. [^64] Note that in the first order in $\epsilon$ it does not matter whether we choose $A^{\prime}=A$ or $A^{\prime}=\widetilde{A}$ in the cohomology comparison map $\mathfrak{B}_{\widetilde{A}\leftarrow A;A^{\prime}}$. The r.h.s. of equation (219) reads
@@
(223) $=[\varphi_{A,A}(\mathsf{a})+\epsilon(d_{\mathsf{a}}\varphi_{A,A})(\alpha+\mathsf{b}+\beta)].$
@@
(226) $\det(B^{\vee})\circ\underline{Z}_{\underline{\varphi}_{A}(\alpha)}(B(\mathsf{a}))=\underline{Z}_{A}(\alpha+\mathsf{a})-i\hbar\Delta_{\mathsf{a}}R(A,\alpha,\mathsf{a}),$
@@
(229) $\nabla^{G}\underline{Z}=-i\hbar\Delta_{\mathsf{a}}(\underline{r}\,\underline{Z}),$
@@
(230) $\nabla^{G,\mathrm{triv}}\underline{Z}=-i\hbar\Delta_{\mathsf{a}}(\pi_{*}I^{*}(r)\,\underline{Z}),$
@@
(256) $Z^{\mathrm{glob}}$ $=\sum_{k=0}^{\dim\mathcal{M}^{\prime}}\frac{(-i\hbar)^{k}}{k!}\left\langle\frac{\partial}{\partial\mathsf{a}^{2}},\frac{\partial}{\partial[\delta A]}\right\rangle^{k}\widetilde{Z}^{(k)}\Big|_{\mathsf{a}^{1}=0}$
$=\left(e^{-i\hbar\left\langle\frac{\partial}{\partial\mathsf{a}^{2}},\frac{\partial}{\partial[\delta A]}\right\rangle}\widetilde{Z}\right)\Big|_{\mathsf{a}^{1}=[\delta A]=0}.$
@@
(258) $\mathsf{K}\omega|_{\mathsf{a}^{1}\rightarrow 0}\sim-\frac{1}{\deg\omega}\left\langle\mathsf{a}^{1},\frac{\partial}{\partial[\delta A]}\right\rangle\omega|_{\mathsf{a}^{1}=0},$
@@
Formula (256) can be seen as the perturbative evaluation of the following path integral: [^68] The sign convention for the Berezinian in the outer integral is: $\mathcal{D}\mathsf{a}^{2}\,\mathcal{D}\zeta=\prod_{a}\mathcal{D}\mathsf{a}^{2}_{a}\,\mathcal{D}\zeta^{a}$, for $(\mathsf{a}^{2}_{a},\zeta^{a})$ coordinates on $H^{2}_{A}[-1]\oplus H^{1}_{A}[1]$ associated to some basis $\chi_{a}$ in $H^{1}_{A}$ and the dual basis $\chi^{a}$ in $H^{2}_{A}$.
(260) $Z^{\mathrm{glob}}(A)=\int_{H^{2}_{A}[-1]\oplus H^{1}_{A}[1]}\mathcal{D}\mathsf{a}^{2}\,\mathcal{D}\zeta\int_{\mathcal{L}=\Omega_{d_{A}^{*}\mathrm{-ex}}[1]}\mathcal{D}\alpha_{\mathrm{fl}}\\ \exp\frac{i}{\hbar}\Bigg(S_{CS}(A+i(\mathsf{a}^{2})+\alpha_{\mathrm{fl}})+\langle\zeta,\mathsf{a}^{2}\rangle\\ +\int_{M}\frac{1}{2}\langle\alpha_{\mathrm{fl}},d_{A}G\mathrm{ad}^{*}_{i(\zeta)}\alpha_{\mathrm{fl}}\rangle+\langle\alpha_{\mathrm{fl}},d_{A}G\mathrm{ad}^{*}_{i(\zeta)}i(\mathsf{a}^{2})\rangle\Bigg).$
@@ 5.5. Metric dependence of the global partition function
(267) $\check{Z}^{\mathrm{glob,ren}}$ $=\sum_{k=0}^{\dim\mathcal{M}^{\prime}}\frac{(-i\hbar)^{k}}{k!}\left\langle\frac{\partial}{\partial\mathsf{a}^{2}},\frac{\partial}{\partial[\delta A]}\right\rangle^{k}\overline{Z}^{\mathrm{ren}(\bullet,k)}\Big|_{\mathsf{a}^{1}=0}$
$=\left(e^{-i\hbar\left\langle\frac{\partial}{\partial\mathsf{a}^{2}},\frac{\partial}{\partial[\delta A]}\right\rangle}\overline{Z}^{\mathrm{ren}}\right)\Big|_{\mathsf{a}^{1}=[\delta A]=0}.$
@@
    The induced differential in (270) is $\sum_{k\geq 0}\mathsf{p}\varkappa(-\mathsf{K}\varkappa)^{k}\mathsf{i}$. Since the image of $\mathsf{i}$ is in zero-forms on $\mathcal{M}^{\prime}$, $\mathsf{K}$ reduces the form degree on $\mathcal{M}^{\prime}$ by one and $\varkappa$ does not change the form degree on $\mathcal{M}^{\prime}$, all terms with $k>0$ vanish and the induced differential is $\mathsf{p}\varkappa\mathsf{i}=\delta_{g}-i\hbar\Delta$.
@@
(272) $\check{Z}^{\mathrm{glob,ren}}(A,\mathsf{a}^{2}_{0})=e^{\frac{i}{\hbar}c(\hbar)\frac{S_{\mathrm{grav}}(g,\phi)}{2\pi}}\int_{H^{2}_{A}[-1]\oplus H^{1}_{A}[1]}\mathcal{D}\mathsf{a}^{2}_{\mathrm{fl}}\,\mathcal{D}\zeta\int_{\mathcal{L}=\Omega_{d_{A}^{*}\mathrm{-ex}}[1]}\mathcal{D}\alpha_{\mathrm{fl}}\\ \exp\frac{i}{\hbar}\Big(S_{CS}(A+i(\mathsf{a}^{2})+\alpha_{\mathrm{fl}})+\langle\zeta,\mathsf{a}^{2}_{\mathrm{fl}}\rangle\\ +\int_{M}\underbrace{\frac{1}{2}\langle\alpha_{\mathrm{fl}},d_{A}G\mathrm{ad}^{*}_{i(\zeta)}\alpha_{\mathrm{fl}}\rangle+\langle\alpha_{\mathrm{fl}},d_{A}G\mathrm{ad}^{*}_{i(\zeta)}i(\mathsf{a}^{2})\rangle}_{-\frac{1}{2}\langle i(\mathsf{a}^{2})+\alpha_{\mathrm{fl}},H_{\delta A^{\prime}=i(\zeta)}(i(\mathsf{a}^{2})+\alpha_{\mathrm{fl}})\rangle}\underbrace{-\frac{1}{2}\langle\alpha_{\mathrm{fl}},\lambda_{\delta g}\alpha_{\mathrm{fl}}\rangle-\langle\alpha_{\mathrm{fl}},\lambda_{\delta g}i(\mathsf{a}^{2})\rangle}_{-\frac{1}{2}\langle i(\mathsf{a}^{2})+\alpha_{\mathrm{fl}},H_{\delta g}(i(\mathsf{a}^{2})+\alpha_{\mathrm{fl}})\rangle}\Big)\Bigg|_{\mathsf{a}^{2}=\mathsf{a}^{2}_{0}+\mathsf{a}^{2}_{\mathrm{fl}}}$
with $H_{\delta A^{\prime}}$ and $H_{\delta g}$ as in (179b), (179c); $\mathsf{a}^{2}_{0}$ is interpreted as a vector in $T^{*}_{A}[-1]\mathcal{M}^{\prime}$.
@@
(273) Zˇglob,ren(1)=eic()Sgrav(g,ϕ)2πeiSCS(A)eπi4ψA,gτA1/2([Uncaptioned image]+
+[Uncaptioned image]+[Uncaptioned image]+
@@
Here the graphical conventions are as in Figure 13; loose half-edges are decorated by $i(\mathsf{a}^{2}_{0})$; black circle vertex is decorated by $H_{\delta g}$. The graphs shown contribute in zeroth order in $\hbar$ and $\cdots$ is of order $\geq 1$ in $\hbar$.
@@ Appendix D Some technical proofs
D.1. Proof of Propositions 2.12 and 2.13
D.1.1. Proof of Proposition 2.12
@@
(333) $(d\widetilde{\kappa}_{A_{0}})_{\delta}=\mathrm{Id}+K_{A_{0}}\mathrm{ad}_{\mathfrak{d}}\colon\Omega^{1}\to\Omega^{1}$
is invertible. In particular, by the triangle inequality this happens when the operator norm of $K_{A_{0}}\mathfrak{d}$ is less than one.
@@
If $d_{A_{0}}\alpha=-\frac{1}{2}[\alpha,\alpha]$, then the first two terms cancel and the latter two terms vanish because $\alpha,\alpha]=-2d_{A_{0}}\alpha$ is exact. Point iv) follows immediately from the definition of $\widetilde{\varphi}_{A_{0}}=A_{0}+\widetilde{\delta}_{A_{0}}$ and points i) and iii). ∎
D.1.2. Proof of Proposition 2.13
Lemma D.1.
Proof.
We know that $\widetilde{\delta}_{A_{0}}$ converges in a neighborhood $U_{1}$ of $0\in\Omega^{1}_{l}$, therefore $\widetilde{\delta}_{A_{0}}$ is defined on $U=U_{1}\cap\operatorname{Im}i_{A_{0}}$. On $\operatorname{Im}i_{A_{0}}$, we know that $\widetilde{\delta}_{A_{0}}\in MC_{A_{0}}$ by Proposition 2.7. Since $\widetilde{\delta}_{A_{0}}(\alpha)=\alpha+K_{A_{0}}(\ldots)$, we have $K_{A_{0}}\widetilde{\delta}_{A_{0}}(\alpha)=0$. Therefore $\widetilde{\delta}_{A_{0}}(U)\subset MC_{A_{0}}\cap\ker K_{A_{0}}$. On the other hand, if $m\in MC_{A_{0}}\cap\ker K_{A_{0}}$, then $d_{A_{0}}\widetilde{\kappa}(m)=K_{A_{0}}\widetilde{\kappa}(m)=0$, i.e $\widetilde{\kappa}_{A_{0}}(m)\in\operatorname{Im}i_{A_{0}}$. For $m$ small enough, we therefore have $\widetilde{\kappa}_{A_{0}}m\in U$, and since we already know that $\widetilde{\delta}_{A_{0}}$ and $\widetilde{\kappa}_{A_{0}}$ are inverse to each other, we conclude the statement. ∎
Corollary D.2.
The restriction of the map $\kappa_{A_{0}}\colon\Omega^{1}\to H^{1}_{A_{0}}$ given by
(334) $\kappa_{A_{0}}(\mathfrak{d})=p_{A_{0}}(\mathfrak{d}+\frac{1}{2}K_{A_{0}}[\mathfrak{d},\mathfrak{d}])=p_{A_{0}}(\mathfrak{d})\in H^{1}_{A_{0}}$
Finally, let us prove Proposition 2.13.
Proof of Proposition 2.13.
We begin with point i). To show surjectivity onto a small neighborhood of $A_{0}\in\mathrm{FC}_{l}$, in the first step, we construct a gauge transformation that takes an arbitrary connection $A_{1}$ close enough to $A_{0}$ to a connection $A_{1}^{\prime}$ satisfying $K_{A_{0}}(A_{1}^{\prime}-A_{0})=0$. To this end, consider the map $F\colon\Omega^{0}_{K-\mathrm{ex},l}\times\Omega^{1}_{l}\to\Omega^{0}_{K-\mathrm{ex},l}$ given by
$(\beta,\mathfrak{d})\to K_{A_{0}}(\,{}^{\exp(-\beta)}(A_{0}+\mathfrak{d})-A_{0}).$
We want to solve for $\beta=\beta(\mathfrak{d})$ such that $F(\beta(\mathfrak{d}),\mathfrak{d})=0$, this is the desired gauge transformation. Existence of $\beta$, for small enough $\mathfrak{d}$, is then guaranteed by the implicit function theorem for Banach spaces, since the derivative of $F$ at $(0,0)$ in direction of $\beta$ is $(dF/d\beta)(0,0)=K_{A_{0}}d_{A_{0}}=\mathrm{id}_{\Omega^{0}_{K-\mathrm{ex}}}$. This means that for $A_{1}$ close enough to $A_{0}$, there is $\beta$ such that $K_{A_{0}}(\,{}^{\exp(-\beta)}A_{1}-A_{0})=0$. For such connections, we know that $\widetilde{\kappa}_{A_{0}}(\,{}^{\exp(-\beta)}A_{1}-A_{0})$ is a $d_{A_{0}}$- and $K_{A_{0}}$-closed 1-form. Therefore, if $A_{1}$ is a flat connection close to $A_{0}$, then $\alpha=\widetilde{\kappa}_{A_{0}}(\,{}^{\exp(-\beta)}A_{1}-A_{0})-d_{A_{0}}\beta$ is a $d_{A_{0}}$-closed form such that $\widetilde{\psi}_{A_{0}}(\alpha)=A_{1}$.
As for point ii), we notice that the implicit function theorem also guarantees smoothness of the map $\mathfrak{d}\mapsto\beta(\mathfrak{d})$. It follows that $\widetilde{\psi}_{A_{1}}^{-1}\circ\widetilde{\psi}_{A_{0}}$ is a composition of smooth maps between differential forms (projections, gauge transformations, $\beta$, and the maps $\widetilde{\varphi}$ and $\widetilde{\kappa}$) and hence smooth. ∎
@@
However, by the results of Axelrod-Singer [AS91], the singular terms of the heat kernel expansion are proportional to $\mathrm{id}\in\mathrm{End}(\mathfrak{g})$ and hence vanish under the trace with $\mathrm{ad}_{\dot{A}_{0}}$ by unimodularity of $\mathfrak{g}$. Therefore, the zeta-regularized supertrace coincides with the point-splitting regularized supertrace that we use to define tadpoles in Feynman diagrams, cf. footnote 45.